What are the methods to be studied in determining atomic weight and molecular weight?

Grade:12th Pass

1 Answers

Arun
5 years ago

Following methods are used to determine molecular mass,

(i) Diffusion method (For gases)The ratio of rates of diffusion of two gases is inversely proportional to the square root of their molecular masses.

r1/r= √m2/m1

(ii) Vapour density method (For gases only): Mass of a fixed volume of the vapour is compared with the mass of the same volume of hydrogen under same conditions. The ratio of these masses is called Vapour density or Relative density.

Molecular mass = 2×Vapour desity   

(iii) Victor Meyer method (For volatile liquids or solids)

It is based on Dalton's law of partial pressure and Avogadro's hypothesis (gram molar volume).

22400 ml of vapours of a substance = Molecular mass of that substance

Average atomic mass and molecular mass

A- (Average atomic mass) = ΣAi Xi/ΣXtotal

M- (Average molecular mass) = ΣMi Xi/ΣXtotal

Where A1 A2, A3…. are atomic mass of species 1, 2, 3,…. etc. with % ratio as XX2 X3…etc. Similar terms are for molecular masses.